When 50 ml water sample titrated against 0.02 N HCl acid using phenolphthalein


indicator, the titration reading was 24.2 ml. After boiling 100 ml of the same sample using


methyl orange indicator, the titration value was 28.2 ml. Calculate permanent and total hardness.

1
Expert's answer
2022-01-17T02:02:01-0500

28.2-24.2= 4

100-50= 50

= 50/4= 12.5